I could really use some help with these questions!

1.

Question 1 Assume a population of 1,000 people of when 15 are sick with a certain disease, 12 of the 15 die from this disease. Calculate the mortality rate from the disease

2.

3.
Answer #1

1) Mortality rate is number of deaths divided by total number of population.

Here, No. of deaths= 12

No. of population= 1000

Mortality rate = 12/ 1000

= 0.012

2) Case fatality rate = No. of deaths/ No. of cases × 100

= 12/15×100

= 80%

3) Five new cases are occured during month of July.

Hence, incidence in July is 5/800

4) The incidence rate= 34/ ( 45000-234)

= 34/ 44766

= 76 per 100,000 population.